OpenStudy (xximhisgirlxx):

Does the set of numbers 13, 21, and 24 form a Pythagorean triple? Explain. yes; 132 + 212 ≠ 242 no; 132 + 212 = 242 no; 132 + 212 ≠ 242 yes; 132 + 212 = 242

Mehek (mehek14):

this is how you find it 13^2+21^2=24^2 if it equals, then it forms a pythagorean triple

Mehek (mehek14):

can you solve that?

OpenStudy (xximhisgirlxx):

hang on let me try

OpenStudy (xximhisgirlxx):

yes you can

Mehek (mehek14):

Ok what is 13*13?

OpenStudy (xximhisgirlxx):

169

Mehek (mehek14):

ok what is 21*21

Mehek (mehek14):

look if you want help, you have to cooperate.

OpenStudy (xximhisgirlxx):

21 * 21 is 441

Mehek (mehek14):

and 24*24

OpenStudy (xximhisgirlxx):

576

Mehek (mehek14):

ok now what does 169 + 441 = ?

OpenStudy (xximhisgirlxx):

610

OpenStudy (xximhisgirlxx):

?

Mehek (mehek14):

ok so does 660 = 576?

OpenStudy (xximhisgirlxx):

no

Mehek (mehek14):

than the answer is either b or c what do you think?

OpenStudy (xximhisgirlxx):

c?

Mehek (mehek14):

yup

OpenStudy (xximhisgirlxx):

thank you! you mind helping with some more?